OpenStudy (anonymous):

find x 6arc sin x-pi=0

OpenStudy (anonymous):

\[6 arc \sin x-\pi=0\]

OpenStudy (anonymous):

help @welshfella !!!

OpenStudy (welshfella):

arcsin x mean ' the angle whose sine is x'

OpenStudy (anonymous):

\[\sin^{-1} x=\frac{ \pi }{ 6}\] I did this but I don't know what is next? @welshfella

OpenStudy (welshfella):

well pi/6 is the answer - the angle is pi/6 radians

OpenStudy (welshfella):

x = pi/6

OpenStudy (anonymous):

thanks ! :P

OpenStudy (welshfella):

yw